OpenStudy (anonymous):

the surface area of a spherical scented candle is 19.623cm^2. you want to stack 4 in a tube. how tall does the tube have to be to the nearest tenth?

OpenStudy (anonymous):

Here's a plan: Use surface area formula to solve for radius; double to get diameter, then multiply by 4 to get total height. Round up.

OpenStudy (anonymous):

your stacking them side by side? or one on top of the other?

OpenStudy (anonymous):

on top of each other

OpenStudy (anonymous):

surface area is 2 circles, and a cylindar pi r^2 + pi r^2 + 2pi r h = Sa, solve for h 2pi r^2+ 2pi r h = Sa 2pi r(r+ h) = Sa r+ h = Sa/2pi r h = Sa/2pi r - r, times 4 H = 4(Sa/2pi r - r)

OpenStudy (anonymous):

you really need to know the radius to determine a concrete value

OpenStudy (anonymous):

or the height of 1 candle

OpenStudy (anonymous):

Ok, I think i got it thank you!

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

Thank you very much!